# Bubble dynamics in a cavitating venturi

### v1

@@ Abstract
Cryogenic fluids have extensive applications as fuel for launch vehicles in space applications and research. The physics of cryogenic flows are highly complex due to the sensitive nature of phase transformation from liquid to bubbly liquid and vapor, eventually resulting in cavitating flows at the ambient temperature owing to the very low boiling point of cryogenic fluids, which asserts us to classify such flows under multi-phase flow physics regime. This work elucidates the modeling of bubbly flow for cryogenic fluids such as liquid nitrogen in a converging-diverging venturi-like flow device known as cavitating venturi, a passive flow control metering device. The numerical works in literature are usually limited to modeling iso-thermal bubbly flows such as water devoid of involving energy equations because there is no occurrence of interface heat transfer as latent heat of vaporization of water is higher, unlike cryogenic fluids which are sensitive to phase change at ambient conditions. So, to realize an appropriate model for modeling cryogenic bubbly flows such as liquid nitrogen flow, the effect of heat transfer at the interface and convective heat transfer from the surrounding liquid to the traversing bubble needs to be included. Numerical modeling using an in-house code involving a finite-difference method The numerical results showed the importance of including the heat transport equation due to convection and at the interface of bubble-fluid as a significant source term for the bubble dynamics. The work is supported by computational simulation using a commercial CFD package for 2-dimensional simulations to predict a characterizing parameter, namely cavitation length. A limited flow visualization experiment using a high-speed camera is performed to study the cavitating zone length.
@@ 1 Introduction
A cavitating venturi is a simple passive flow and metering device that ensures a constant mass flow rate of the fluid that passes through it, even under varying downstream pressure conditions. The cavitating venturi has a converging section followed by a throat and a diverging section. In a venturi, the throat has a minimum area. Because of this reason, whenever a flow happens in a cavitating Venturi, the fluid suffers from minimum pressure near the throat. It so happens that the fluid pressure attains the vapor pressure, which results in cavities or vapor bubbles forming, and the flow begins to cavitate. Mixed behavior, such as bubble formation and bubble collapse from the throat to the diverging section, is typical of cavitating flows in the venturi. Such a two-phase flow is also identified with nucleation, bubble growth and collapse, fission, fusion of bubbles, re-entrant jet formation, and turbulent mixing of vapor bubbles.
Cavitating flow is a typical two-phase flow phenomenon that is common in mechanical systems like Turbo pumps (near inducer), Marine propellers, etc. When pumps operate at high rotational speeds, it results in minimum pressure near the inducer, which in turn causes vapor formation resulting in attached cavitation and traveling bubble cavitation, which are the usual cavitation patterns commonly observed in turbomachines. cavitation causes considerable damages like erosion on the surface of the impeller vanes of the pumps and eventually reduces its service life.
Besides serious disadvantages due to cavitation, cavitation positively impacts a passive flow device like cavitating Venturi, which is discussed in the following section. The following figure 2 shows cavitation in a typical cavitating Venturi.
@@
Cavitating Venturi is a Passive flow control and metering device that does not require any complicated flow control instrumentation like a Controlled valve and the associated control systems. Instead, a passive flow device is a simple mechanical device that works fundamentally on the critical or chocked flow principle. The following figure 1 depicts the Rocket propulsion system with active and passive fuel flow control systems.
@@
Both active and passive flow control systems are commonly used in Rocket propulsion flow systems to supply liquid propellants from the fuel tank to the combustion chamber via fuel injectors. However, active flow control systems are complicated by signal actuating mechanisms for control valves, owing to which they lose their reliability. Instead, a Passive flow system equipped with a Cavitating Venturi in place of a flow control valve seems advantageous in ensuring the reliability of the flow system, owing to its simple operating mechanism, namely choked flow.
Cavitation plays a positive role in Cavitating Venturi used in Passive Flow Control systems, as Cavitation helps maintain a constant flow rate. This Constant flow rate is achieved because of critical flow or chocked flow phenomenon.
  • The critical flow condition is common in both Compressible &\&& Incompressible flows. In a Compressible flow, it is understood that the flow becomes simply chocked by accelerating the fluid to the local sonic velocity at the throat section of the nozzle (as downstream pressure waves cannot travel upstream).
  • For an incompressible fluid, the sonic velocity is significantly greater than that of a compressible fluid. So, in a Nozzle or Venturi, the accelerating incompressible fluid starts evaporating before reaching its sonic velocity, and vapor bubbles appear near the throat section and further downstream.
  • The so-formed vapor bubble cloud volume is large, obstructing the fluid flow at the throat. As pressure decreases, cavitation intensity increases, which causes the area of the bubble cloud to increase further. As a result, there is an increased resistance to the flow near the throat location, which ensures the flow rate remains unchanged, exhibiting a Cavitating flow condition.
  • The formed Vapour bubbles collapse back to ensure a single-phase liquid flow in the diverging section of the Cavitating Venturi because of the gradual increase in fluid Pressure along the diverging section.
Because of this attractive property, Cavitating Venturi finds application in Liquid and Cryogenic Stages of Rocket Engines, as fluctuating pressures in the Combustion chamber, which is equal to Venturi downstream pressure, will not affect the quantity of flow through Cavitating Venturi for a wide range of operating conditions. Much research has been done to develop numerical models to predict the characteristics of choked flow[2-4,9-11]. Also, many experimental works[5,7-8] have been aimed at understanding the phenomenon and behaviour of cavitation. It also appears from open literature that very few works exist in cryogenic Cavitating Venturi flows.
The behaviour of Cryogenic cavitating flows in a Venturi is very complex because of the very low Latent heat of Vaporization of the cryogenic fluid and hence requires a detailed study of flow behavior experimentally and validating the same using an appropriate mathematical model which can capture the actual physics of Cavitating flow behavior in a Venturi.
The following are the brief of the objectives put forth
  1. 1.
    Systematically study the applicability of 1D models and the predictive capabilities of a typical one-dimensional model that could be used for Cryogenic cavitating venturi. Though 1D models are not very rigorous in ultimately predicting the behavior of cavitating venturi, they could be very useful in predicting some key parameters. There have been many studies in 1D models and are gradually being improved for cavitating venturi with water as the working fluid. However, there seems to be limited published work on extending the numerical model and trying to characterize the behaviour of cryogenic cavitating venturi. Owing to the limitations in the 1D model, it naturally leads to the following second objective.
  2. 2.
    To explore the predictive capabilities of a 2D model of cryogenic cavitating venturi using commercial CFD solver ANSYS fluent, incorporating appropriate cavitation model. These are expected to give realistic estimates of the nature of cavitation and the cavitation length for a given set of operating conditions. Such models, however, require tuning with the experimental results. Hence, the third objective of experimental work was decided.
  3. 3.
@@
    To design and develop an experimental facility with a planar venturi test section incorporating appropriate visualization features to understand flow behaviour when handling cryogenic liquid Nitrogen and obtain the cavitating length.
The work of Brennen [9] has been taken as a basis and appropriately extended with the inclusion of thermal term for cryogenic fluids. The numerical methodology for cavitation prediction has been extended further with 2D simulations using Ansys Fluent.
@@ 2 Numerical 1D model
In Brennen [9] model, the Rayleigh Plesset Equation for the isothermal case was assumed for solving the bubble dynamics. However, this bubble dynamics equation cannot be extended to the case of cryogenic liquids, as heat transfer effects are predominant in cryogenic liquids. It is advantageous to have an additional term in the same Brennen model [9] that accounts for heat transfer and extends for cryogenic cavitating venturi flows. The following section details models used for Numerical 1D studies for cryogenic cavitating flows.
@@ 2.1 Derivation of modified Rayleigh-Plesset equation
The above equation is the simple Rayleigh-Plesset Equation for the isothermal case. The equation does not have an appropriate term for handling the heat interactions. In the above equation the vapour pressure inside the bubble which can be represented as $P_{v}(T_{B})$. In ordinary fluid such as water, there is no temperature difference between the bubble temperature $T_{B}$ and the surrounding liquid temperature $T_{\infty}$ because of the isothermal conditions. But this assumption is not valid for cryogenic liquids as the slightest temperature difference between $T_{B}$ and $T_{\infty}$ causes a considerable change in the vapour pressure inside the bubble. If there is a difference in temperature between the bubble and the surrounding liquid, then the temperature gradient $T_{\infty}-T_{B}$ becomes a potential source for heat transfer influencing the bubble growth.
@@
The third term on the right-hand side of the equation is referred to as a thermal term, and the second term is the tension or the driving term for bubble growth. To find an expression for the thermal term, we need to look into the Heat Balance and B-factor theory [1] & [5]
The thermal effect in cryogenic liquids can be interpreted as the heat interaction between the vapour bubbles and the surrounding liquid medium. The production of a vapour volume $V_{\nu}$ requires the quantity of heat $\rho_{\nu}V_{\nu}L$ to be supplied (L is the latent heat of vaporization). This energy is given by the liquid near the vapour cavity. So, the liquid mass surrounding the vapour bubble gets cooled by supplying heat to the bubble. Steffano [1] & [5] put forth a theory, popularly known as B-factor theory, which gives a non-dimensional number for temperature drop called B-factor [1]
@@
From the above equation, the B-factor is given by the ratio of the volume of vapor to the volume of liquid. Alternately, it is also given by the ratio of the actual temperature drop to the reference temperature drop, which is shown below
@@
This slight temperature drop $\Delta T$ in the liquid happens to be the reason for the difference in temperature between a vapor bubble and the liquid in the immediate vicinity, which affects the vapor pressure change inside the bubble, as the bubble temperature has now been slightly changed. If the drop in temperature is very small, then by Taylor series expansion, we can write
@@
The vapour pressure change with respect to the temperature $\dfrac{dP_{\nu}}{dT}$ is given by Clausius Clayperon relation[8]
@@
To estimate the temperature depression $\Delta T$, heat transfer between the liquid and bubble has to be considered. In literature, both conductive and convective modes of heat transfer were assumed based on the flow conditions, i.e., if there is no slip between liquid and bubble, then the conductive approach is sufficient. But in our case, assuming a convective mode of heat transfer is correct.
Assuming $Q$ as the heat flux which is exchanged at the interface of bubble and liquid, then we can write
@@
Positive $\delta T$ means temperature decrement in liquid and temperature increment in vapour bubble. This, in turn, increases the vapour pressure inside the bubble, which in turn makes the bubble expand or grow. Similarly, it can be interpreted that negative $\delta T$ leads to shrinkage of the vapor bubble due to condensation.
@@ 2.2 Modelling 1D cavitating nozzle / Venturi flow
Similar to Wang and Brennen Quasi-1D model [9] for Bubbly Cavitating flows through a Nozzle, a Quasi-1D model with thermal effects is attempted in this work for applicability with cryogenic liquids. The Continuity and momentum equations for bubbly two-phase flows along with closures, i.e., the volume fraction equation and the modified Rayleigh Plesset Equation along with the Bubble momentum equation for estimation of bubble velocity, were considered in this model for solving bubble radius
For modelling the heat transfer between bubble and liquid, a convective heat transfer model was taken from the existing Ranz and Marshall model [14]. The turbulence model was not incorporated in 1D numerical studies as handling such equations was tedious. Following assumptions were made in this model,
@@
    At any cross-section of the Venturi, the bubbles had a uniform size, and their number was assumed constant.
@@
Supposing that flow develops only in the x direction, the steady bubbly flow continuity and momentum equations that Wang and Brennen used are given as follows
@@
Where, $C_{p}(x,t)=\dfrac{(P(x,t)-P_{0})}{(0.5\rho_{l}u_{0}^{2})}$ is the fluid pressure coefficient $P(x,t)$ is the fluid pressure , $P_{0}$ is the upstream fluid pressure, $u_{0}$ is the upstream fluid velocity and $u$ is the velocity along the length of the profile. $A(x)$ corresponds to the area profile, which is the function of space coordinate, i.e., $x$ (being a 1D case). The mathematical expression for the area profile A(x) is
@@
$\alpha(x,t)=\dfrac{\dfrac{4}{3}\pi\eta R(x,t)^{3}}{\left(1+\dfrac{4}{3}\pi\eta R(x,t)^{3}\right)}$ is the non-dimensional void fraction, $R(x,t)$ is the non-dimensional bubble radius and $\eta$ is the non-dimensional population number.
@@
In the above equation, the last term on the right hand side is the thermal term, and $h_{b}$ is the convective heat transfer coefficient which was modelled using the already existing Ranz and Marshall [14] model for the study of evaporation of droplets in spray drying which is given by
@@
Where $Nu_{b}=2+0.6Reb^{\frac{1}{2}}Pr^{\frac{1}{3}}$ is the bubble Nusselt number, which in turn depends on bubble Reynolds number ($Re_{b}$), and Prandtl number (Pr) which is given by [8] as
@@ 2.3 Wang and Brennen benchmark Results for Water with thermal term
A numerical 1D code using Matlab was written for solving the Brennen model for water as the fluid with the effects of the thermal term. The results were plotted for bubble radius, void fraction, coefficient of pressure, velocity of the liquid phase, and the growth contribution terms (thermal and pressure terms). The non-dimensional bubble radius, coefficient of pressure, void fraction, and non-dimensional void fraction plots are shown below in the figure 3,7, 8, 9corresponds to the Rayleigh Plesset bubble dynamics equation with the thermal effects incorporated.
@@
The figure 10 gives the plot of source terms in the Rayleigh-Plesset equation. The thermal source term is compared with a pressure source term. The values of the same are plotted for an initial void fraction of $\alpha_{0}=2.5\times 10^{-6}$, where the term $\alpha_{0}$ is the initial void fraction which has been arbitrarily assumed in Wang and Brennen model [9]. This initial void fraction significantly impacts the numerical prediction of bubble radius as this signifies the number of nucleation sites or microscopic voids in the liquid phase.
For getting a clear understanding of the domination of different source terms in the bubble dynamics, the right-hand side of the modified Rayleigh-Plesset equation containing such source terms was handled separately, and the magnitude of the same was plotted. i.e., all other terms except the thermal term were taken as the Pressure source term. The following figure 10,11 shows the contribution of the above-mentioned thermal term and pressure term.
@@
It can be inferred from the plots that the contribution of the thermal term within the flow domain of the venturi is small and is of the order $10^{-1}$ against the contribution from the pressure term, which is of the order $10^{4}$. This points out to the fact that the explicit modeling of the thermal term in the modified Rayleigh Plesset Equation may not provide pronounced effects for fluids with the high latent heat of vapourization. This fact explains the close prediction of the Wang &\&& Brennen model results [9], where such a thermal term is not modeled.
@@
However, the present one-dimensional model is far from providing estimates for performance prediction of cavitating venturi, and more systematic work in developing proper mathematical models to predict the performance would be required. Hence, a systematic study of a simple planar 2D numerical study was attempted using ANSYS fluent, detailed in the following section.

### v2

@@ Abstract
Cryogenic fluids are widely used as fuel for launch vehicles and many other industrial applications. The physics of cryogenic flows is highly complex due to their sensitivity to phase changes. Because of their low boiling points, cryogenic liquids undergo phase transition from liquid to bubbly liquid to vapour. This leads to cavitating flows that fall under the two-phase flow regime. This study models bubbly flows of cryogenic fluids, specifically liquid nitrogen, in a converging-diverging venturi device known as a cavitating venturi, a passive flow control and measurement device. Existing numerical studies in the literature are mostly limited to isothermal bubbly flows, such as water, where energy equations are often neglected due to minimal heat transfer at the interface. In contrast, cryogenic fluids are highly sensitive to phase changes under ambient conditions and interface heat transfer plays a significant role owing to their low latent heat of vaporization. To develop an appropriate model for cryogenic bubbly flows, we incorporate the effects of convective heat transfer from the quiescent liquid to the traversing bubble and heat transfer at the bubble-fluid interface. The numerical modelling is carried out using an in-house finite-difference code, and the results highlight the critical influence of heat transport equations on bubble dynamics. A commercial CFD package is used for two-dimensional simulations to complement the in-house simulations to predict cavitation length, a key characterizing parameter. In addition, limited flow visualization experiments are performed using a high-speed camera to analyze the length of the cavitation zone. This combined numerical and experimental approach provides valuable insight into the dynamics of cryogenic bubbly flows in cavitating venturi systems.
@@ 1 Introduction
A cavitating venturi is a simple, passive flow control and metering device that maintains a constant mass flow rate of fluid, even when downstream pressure conditions fluctuate. The venturi consists of three key sections: a converging section, a throat, and a diverging section. The throat, having the minimum cross-sectional area, causes the fluid pressure to drop to its lowest point as the flow passes through it. When the fluid pressure near the throat reaches its vapour pressure, vapour cavities or bubbles begin to form, initiating cavitation [1]. In a cavitating venturi, the flow exhibits a mix of behaviours, including bubble formation, growth, collapse, and dynamic interactions such as fission, fusion, re-entrant jet formation, and turbulent mixing of vapour bubbles as the flow moves from the throat into the diverging section. These characteristics are typical of cavitating two-phase flows. Lord Rayleigh’s [2] work was one of the earlier works on cavitation dynamics. Plesset et al. [3] put forth a modified bubble dynamics equation for the present numerical work. Many theoretical, numerical, and experimental verification literature exists on cavitating flows and in particular the cavitating flows in converging diverging nozzles or cavitating venturi [4]-[10]. Cavitating flow is a well-known two-phase phenomenon commonly encountered in mechanical systems such as turbo pumps (near the inducer) and marine propellers. In pumps operating at high rotational speeds, the pressure near the inducer can drop to levels that induce vapour formation, resulting in cavitation patterns such as attached cavitation and travelling bubble cavitation. These cavitation patterns are often observed in other turbomachines and can cause significant damage, including erosion of impeller vane surfaces, which ultimately shortens the service life. Despite its detrimental effects on many mechanical systems, cavitation can benefit passive flow control devices such as cavitating venturi. The figure below (Figure 2) illustrates the development of cavitation in a typical cavitating venturi.
@@
As a passive flow measurement device, a cavitating venturi eliminates the need for complex control valves and actuating systems. Instead, it leverages the choked flow principle to maintain a constant flow rate, regardless of downstream pressure variations. Figure 2 illustrates developed cavitation in a venturi.
@@
Rocket propulsion systems rely on flow control to deliver liquid propellants to combustion chambers. While active systems involve control valves and actuators, their reliability is compromised by mechanical complexity. In contrast, a cavitating venturi serves as a reliable passive alternative, using cavitation-induced flow restriction to stabilize the flow rate.
In a cavitating venturi:
(a). For compressible flows, choked flow occurs as fluid accelerates to sonic velocity at the throat, preventing upstream pressure propagation.
(b). For incompressible flows, evaporation occurs before sonic velocity is reached, forming vapor bubbles near the throat. These bubbles restrict the flow, ensuring a stable flow rate.
(c). Downstream, pressure recovery collapses the vapor bubbles, restoring single-phase flow.
This unique property makes cavitating venturi ideal for liquid and cryogenic rocket engines, where flow remains steady despite fluctuations in combustion chamber pressure. Cryogenic cavitating flows, due to their low latent heat of vaporization, exhibit complex behavior which mandates a detailed flow physics study. The key objectives of this work are:
1. To assess the predictive capabilities of one-dimensional models for cryogenic cavitating venturis. Though simplified, 1D models provide valuable estimates of key flow parameters.
2. To develop two-dimensional models using ANSYS Fluent with appropriate cavitation models. Also to predict flow characteristics, such as cavitation length, and validate against experiments.
@@
3. Design and develop a planar venturi test facility with flow visualization features. Conduct experiments using cryogenic liquid nitrogen to measure cavitation length and flow behavior.
This study builds upon Wang and Brennen’s approach [9], extending it to include thermal effects in cryogenic flows. Numerical simulations with 2D models provide a robust framework to capture the complex dynamics of cavitating venturi systems.
@@ 2 Numerical 1D model
In Brennen [9] model, the Rayleigh Plesset Equation for the isothermal case was assumed for solving the bubble dynamics. However, this bubble dynamics equation cannot be extended to the case of cryogenic liquids, as heat transfer effects are predominant in cryogenic liquids. It is advantageous to have an additional term in the same Brennen model [9] that accounts for heat transfer and extends to cryogenic cavitating venturi flows. The following section details models used for the Numerical 1D studies for cryogenic cavitating flows.
@@ 2.1 Derivation of modified Rayleigh-Plesset equation
The above is the simple Rayleigh-Plesset equation for the isothermal case. The equation does not have an appropriate term for handling the heat interactions. The vapour pressure inside the bubble can be represented as $P_{v}(T_{B})$ in this equation. In an iso-thermal fluid such as water, there is no temperature difference between the bubble temperature $T_{B}$ and the surrounding liquid temperature $T_{\infty}$. The above isothermal condition is an assumption which is not valid for cryogenic liquids as the slightest temperature difference between $T_{B}$ and $T_{\infty}$ causes a considerable change in the vapour pressure inside the bubble. The temperature gradient $T_{\infty}-T_{B}$ becomes a potential source for heat transfer influencing the bubble growth.
@@
The third term on the right-hand side of the equation is the thermal term, and the second term is the driving term for the bubble growth. To find an expression for the thermal term, we need to look into the heat Balance and B-factor theory [1] & [5]
The thermal effect in cryogenic liquids can be interpreted as the heat interaction between the vapour bubbles and the surrounding liquid. The production of a vapour volume $V_{\nu}$ requires the quantity of heat $\rho_{\nu}V_{\nu}L$ to be supplied (L is the latent heat of vaporization). This energy is given by the liquid near the vapour cavity. Hence, the liquid mass surrounding the vapour bubble gets cooled by supplying heat to the bubble. Steffano [1] & [5] put forth a theory, popularly known as B-factor theory, which gives a non-dimensional number for temperature drop called B-factor [1]
@@
From the above equation, the B-factor gives the ratio of the volume of vapour to that of the volume of liquid. Alternately, it is also given by the ratio of the actual temperature drop to the reference temperature drop, which is shown below
@@
This slight temperature drop $\Delta T$ in the liquid happens to be the reason for the difference in temperature between a vapour bubble and the liquid in the immediate vicinity, which affects the vapor pressure change inside the bubble, as the bubble temperature has now been slightly changed. If the drop in temperature is very small, then by Taylor series expansion, we can write
@@
The vapour pressure changes with respect to the temperature $\dfrac{dP_{\nu}}{dT}$ which is given by the Clausius Clayperon relation[8]
@@
To estimate the temperature difference $\Delta T$, the heat transfer between liquid and vapour bubble has to be quantified. In literature, both conductive and convective modes of heat transfer were assumed based on the flow conditions, i.e., if there is no slip between liquid and bubble, then the conductive approach is sufficient. But in our case, assuming a convective mode of heat transfer is applicable.
Assuming $Q$ as the heat flux at the interface of bubble and liquid, then we can write
@@
Positive $\delta T$ indicates a temperature decrement in liquid, but a temperature increment in vapour bubble. This, in turn, increases the vapour pressure inside the bubble, which in turn makes the bubble expand or grow. Similarly, it can be interpreted that the negative $\delta T$ leads to shrinkage of the vapour bubble due to condensation.
@@ 2.2 Modelling 1D cavitating nozzle / Venturi flow
Similar to Wang and Brennen Quasi-1D model [9] for Bubbly Cavitating flows through a nozzle, a Quasi-1D model with thermal effects is attempted in this work applicable to cryogenic liquids. The continuity and momentum equations for bubbly two-phase flows along with the closures, i.e., the volume fraction equation and the modified Rayleigh Plesset equation along with the bubble momentum equation for estimation of bubble velocity, were considered in this model for solving bubble radius.
For modelling the heat transfer between bubble and liquid, a convective heat transfer model was taken from the existing Ranz and Marshall model [14]. The turbulence model was not incorporated in 1D numerical studies, as handling such equations would be tedious. Following assumptions were made in this model,
@@
    At any cross-section of the venturi, the bubbles had a uniform size, and the bubble number was assumed constant.
@@
Supposing that flow develops only in the longitudinal x direction, the steady bubbly flow continuity and momentum equations that Wang and Brennen used are given as follows
@@
Where, $C_{p}(x,t)=\dfrac{(P(x,t)-P_{0})}{(0.5\rho_{l}u_{0}^{2})}$ is the fluid pressure coefficient $P(x,t)$ is the fluid pressure , $P_{0}$ is the upstream fluid pressure, $u_{0}$ is the upstream fluid velocity, and $u$ is the velocity along the length of the profile. $A(x)$ corresponds to the area profile, which is the function of $x$ (being a 1D case). The mathematical expression for the area profile A(x) is
@@
$\alpha(x,t)=\dfrac{\dfrac{4}{3}\pi\eta R(x,t)^{3}}{\left(1+\dfrac{4}{3}\pi\eta R(x,t)^{3}\right)}$ is the non-dimensional void fraction, $R(x,t)$ is the non-dimensional bubble radius, and $\eta$ is the non-dimensional population number.
@@
In the above equation, the last term on the right side is the thermal term, and $h_{b}$ is the convective heat transfer coefficient, which was modelled using the already existing Ranz and Marshall [14] model for the study of evaporation of droplets in spray drying which is given by
@@
Where $Nu_{b}=2+0.6Reb^{\frac{1}{2}}Pr^{\frac{1}{3}}$ is the bubble Nusselt number, which in turn depends on the bubble Reynolds number ($Re_{b}$), and Prandtl number (Pr) which is given by [8] as
@@ 2.3 Wang and Brennen benchmark Results for Water with thermal term
A numerical 1D code using Matlab was written for solving the Brennen model with water as the fluid considering the effects of the thermal term. The results were plotted for the bubble radius, void fraction, coefficient of pressure, velocity of the liquid phase, and the growth contribution terms (thermal and pressure terms). The non-dimensional bubble radius, coefficient of pressure, void fraction, and non-dimensional void fraction plots are shown below in the figure 3,7, 8, 9corresponds to the Rayleigh Plesset bubble dynamics equation with the thermal effects incorporated.
@@
The figure 10 gives the plot of source terms in the Rayleigh-Plesset equation. The thermal source term is compared with a pressure source term. The values of the same are plotted for an initial void fraction of $\alpha_{0}=2.5\times 10^{-6}$, where the term $\alpha_{0}$ is the initial void fraction which has been arbitrarily assumed in the Wang and Brennen model [9]. This initial void fraction significantly impacts the numerical prediction of bubble radius as this signifies the number of nucleation sites or microscopic voids in the liquid phase.
For understanding the domination of different source terms in the bubble dynamics, the right-hand side of the modified Rayleigh-Plesset equation containing such source terms was handled separately, and the magnitude of the same were plotted. i.e., all other terms except the thermal term were taken as the pressure source term. The following figure 10,11 shows the contribution of the above-mentioned thermal term and pressure term.
@@
It can be inferred from the plots that the contribution of the thermal term within the flow domain of the venturi is small and is of the order $10^{-1}$ when compared to the contribution from the pressure term, which is of the order $10^{4}$. This points out to the fact that the explicit modelling of the thermal term in the modified Rayleigh Plesset Equation may not provide pronounced effects for fluids with the high latent heat of vapourization. This fact explains the close prediction of the Wang &\&& Brennen model results , where such a thermal term is not modelled.
@@
However, the present one-dimensional model is far from providing estimates for performance prediction of cavitating venturi, and more systematic work in developing proper mathematical models to predict the performance would be required. Hence, a systematic study of a simple planar 2D numerical study was attempted using ANSYS fluent, which will be discussed.